The comparative temporary investments and inventory balances of a company follow.

Current Year Previous Year
Accounts payable $94,248 $79,200
Long-term debt   46,324   62,600

Based on this information, what is the amount and percentage of increase or decrease that would be shown on a balance sheet with horizontal analysis?

Amount of Change Increase/Decrease Percentage
Accounts payable $ %
Long-term debt $ %
Answer #1
Amount of change Increase/Decrease Percentage
Accounts payable $15,248 Increase 19%
Long term debt $16,276 Decrease 26%

Accounts payable = 94,248 - 79,200 = 15,248

= (15248 ÷ 79200) * 100

= 19%

Long term debt = 46324 - 62600 = $16,276

= (16276 ÷ 62600) * 100

= 26%